3. The DNA packing ratio is one of important parameters in biology. It is defined as voume of genome (size of genome * volume of a DNA base pair) volume of compartment such as nucleus pack The volume of a DNA base pair is Inn/bp (15 points). (a) What is DNA packing ratio of E. Coli (5 points)? Please look at the table of estimate number of biological systems in our canvas. (b) What is DNA packing ratio of yeast (5 points)? The volume of nucleus of yeast is 4 x (c) What is DNA packing ratio of a Lambda phage with 48,502 base pairs in its head with 9 x 10° μ㎡ (5 points)?

media%2F4f1%2F4f1efa47-23ee-4994-9065-a8

media%2F1da%2F1dab0fed-d72d-433b-9972-f8
Answer #1

E.coli:

Size of the genome 5 x 106 bp

Volume of a base pair: 1 nm3 / bp

Volume of genome: 5 x 106 nm3

Cell volume: 1 µm3 = 1000 nm3

P= (5 x 106 nm3)/ 103 nm3 = 5 x 103

Yeast

Size of the genome 12.1 x 106 bp

Volume of a base pair: 1 nm3 / bp

Volume of genome: 12.1 x 106 nm3

Volume of nucleus: 4 x 109 nm3

P= (12.1 x 106 nm3)/ 4 x 109 nm3 = 3.025 x 10-3

Lambda phage

Size of the genome 48,502 bp

Volume of a base pair: 1 nm3/bp

Volume of genome: 48,502 nm3

Volume of head: 9 x 10-5 µm3 = 9 x 10-2 nm3

P= (48,502 nm3)/ 9 x 10-2 nm3 = 538911.11